OpenStudy (nuccioreggie):

What is the simplified expression for 6 to the power of 4 multiplied by 6 to the power of 3 over 6 to the power of 5? 61 62 63 67

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

hint: 6^(4+3-5)=...

OpenStudy (nuccioreggie):

B

OpenStudy (nuccioreggie):

12

OpenStudy (michele_laino):

hint: 6^(4+3-5)=6^2= 6*6=...?

OpenStudy (michele_laino):

@nuccioreggie

OpenStudy (nuccioreggie):

Ok im goin to solve it right now

OpenStudy (nuccioreggie):

I got 12

OpenStudy (nuccioreggie):

I think its C

OpenStudy (michele_laino):

I got: 6*6=36 so it is B

OpenStudy (nuccioreggie):

So its B 6with the smll 2 on top?